The density of seawater is 1030kg per cubic meter. The pressure exerted in the depth H is 201880Pa. What is the depth of seawater?Take g=9.8N/kg.
1
Expert's answer
2020-08-12T16:45:14-0400

The pressure due to a liquid in liquid columns of constant density is represented by the following formula:


"p = \rho gh"

where "p = 201880\space Pa", "\rho = 1030 \space kg/m^3" is the density of seawater, "g = 9.8 \space N/kg" and "h" is the depth. Expressing "h" obtain:


"h = \dfrac{p}{\rho g} = \dfrac{201880}{1030 \cdot 9.8} = 20 m"

Answer. 20 m.